USE THE FOLLOWING INFORMATION TO ANSWER THE NEXT (3) QUESTIONS:

Mini Corporation factored, with recourse, $600,000 of accounts receivable with Huskie Financing.  The agreement met all three conditions to be considered an outright sale.  Huskie advanced 92% of the amount factored and retained the remainder to cover a 3% finance fee and 5% to cover any sales returns/allowances/discounts.  The recourse obligation is estimated to be 2.4% of accounts factored.

Determine the effect of this transaction on Mini’s financial position: (Use I for increased; D for decreased; or NE for No Effect.  If there is an Effect, state the dollar amount.  Indicate the letter first, then the number. Do not space between the letter and number.  Do not use commas.  For example, if your answer is “Decreased by $4,000”, enter D4000).

1) Determine the effect on Assets

2) Determine the effect on Total Liabilitites

3) Determine the effect on Total Stockholder's Equity

(I saw this same question posted on the website, but all the previous answers appeared to be incorrect.)
Answer #1

Answer:

Important Point keep it Mind:

Mini Corporation factored, with recourse, $600,000 of accounts receivable with Huskie Financing.

Advanced Amount of Factored= 92% i. e ( $600000* 92%) = $552000

Finance Fee= 3% i,e ( $600000* 3%) = $18000

Sales returns/allowances/discounts to cover 5%= ( $600000* 5%) = $30000

The recourse obligation is estimated to be 2.4% of accounts factored. Amount of Recourse Obligation is $600000* 2.4%= $14400.

1) Determine the effect on Assets= D18000. The value of account receivable is $600000. recover upto $582000

2) Determine the effect on Total Liabilitites. Increase by $18000

3) Determine the effect on Total Stockholder's Equity. No effect on Stockholders equity.